OpenStudy (anonymous):

32 is 40 percent of what number? A.12.8 B.128 C.80 D.800 And explanation why?

OpenStudy (anonymous):

C. 80. Explanation: you chance 40% into .4 so the equation is .4 x 80=32

OpenStudy (anonymous):

1) 32 = 40% so 1% = 0.8 (32 divide by 40) so 100% = 80 (x 100) The answer is C.

OpenStudy (anonymous):

32 is 40% of 80; 32 / 0.40 = 80

OpenStudy (jdoe0001):

if 32 is 40% of "n", so "n" is 100% what's "n"? \(\large \begin{array}{cclll} quantity&percentage\\ 32&40\\ n&100\\ \hline\\ \cfrac{32}{n}= \cfrac{40}{100} \end{array}\) solve for "n"